it is reported that 72% of the working women use computers at work. Choose 5 women at random and find the probability that at least 1 doesn't use a computer at work


1
Expert's answer
2022-02-01T17:03:47-0500

Let the random variable "X" be the number of women who do not use computer at work. "X" follows a Binomial distribution with parameters, "n=5" and "p=1-0.72=0.28". That is, "X\sim Bin(n=5,p=0.28)" and it is given as,

"p(X=x)=\binom{5}{x}0.28^x(1-0.28)^{5-x},\space x=0,1,2,3,4,5"


Probability that at least 1woman does not use a computer at work is given as,

"p(x\geq1)=\displaystyle\sum^5_1p(X=x)". To make it simpler, this probability is equivalent to,

"p(x\geq1)=1-p(X=0)"

So,

"p(X=0)=\binom{5}{0}0.28^0(0.72)^5=0.72^5=0.1934918(4dp)".

Thus,

"p(x\geq1)=1-p(X=0)=1-0.1934918= 0.8065082"

Therefore, the probability that at least 1 woman does not use a computer at work is  0.8065082.